OpenStudy (anonymous):

my question is : is sin^2 x , the same as sin x^2?

OpenStudy (anonymous):

No. One is: (sin(x))^2. The other is sin(x^2). Which is not true in general.

OpenStudy (anonymous):

\[sinx ^{2} and (sinx)^{2}\] are different :)

OpenStudy (anonymous):

thanks, whats s the difference though?

OpenStudy (anonymous):

In one, you are evaluating x and THEN squaring it. In the other, you are squaring x THEN evaluating. So for example. Let x=pi/2. For sin(x^2)=sin((pi/2)^2)=sin(pi^2/4)=0.04305 For (sin(x))^2=(sin(pi/2))^2=(1)^2=1. Quite a considerable difference.

OpenStudy (anonymous):

? Where do you get a 2x?

OpenStudy (anonymous):

oh lol read her problem wrong, i saw sin(2x) and sin(x^2) nvm, just listen to malev :D

OpenStudy (anonymous):

Haha :P

OpenStudy (anonymous):

no sin^2(x)